The difference of a number and 6 is the same as 5 times the sum of the number and 2. What is the number


1
Expert's answer
2021-09-22T05:22:18-0400

Let the number be x

The difference of a number and 6 ===> x-6

5 times the sum of the number and 2 ===>5(x+2)

We then have;

x-6=5(x+2)

x-6=5x+10

x-5x=10+6

-4x=16

x=-4

The number is -4